### v1

@@ 1 Introduction
The purpose of this paper is to present sharp estimates for the extremal eigenvalues of the Schur complements arising in saddle point problems of the form (1). If AA is SPD, then the Schur complement $S=BA^{-1}B^{t}$ has the structure of an auxiliary space preconditioner [29, 40], where $A^{-1}$ corresponds to an iterative method defined on the auxiliary space $V$, and $B$ corresponds to the transfer operator mapping the auxiliary space $V$ to the original space $W$. A similar argument also applies in the case when AA is semi-SPD. Thus, the Schur complement $S$ can be analyzed by means of the recently proposed auxiliary space theory [33], which provides sharp estimates for iterative methods for both SPD and semi-SPD systems in a unified framework.
Estimates for the extremal eigenvalues of the Schur complement when AA is SPD are presented in Theorem 4.1, which are sharp in the sense that they are expressed as identities. Sharp estimates for the Schur complement preconditioned by a general SPD preconditioner are given in Theorem 4.5. These results refine existing abstract estimates for the Schur complement (see, e.g., [3, 25, 37]) in the sense that the earlier results can be recovered as straightforward corollaries. While our estimates follow directly from elementary linear algebra within the auxiliary space framework, the existing results required more involved arguments. Finally, our results extend naturally to the case when AA is semi-SPD, as shown in Theorems 4.9 and 4.10.
@@ 1.1 Notation
We summarize here the notation used throughout the paper. Let $V$ and $W$ be finite-dimensional vector spaces equipped with inner products $(\cdot,\cdot)$ and the associated norms $\|\cdot\|$. For any subspace $S\subset V$, we denote by $S^{\perp}$ its orthogonal complement. For a linear operator $B\colon V\to W$, the null space and the range are denoted by $\mathcal{N}(B)$ and $\mathcal{R}(B)$, respectively, and the adjoint of $B$ is denoted by $B^{t}\colon W\to V$.
@@ 2 Auxiliary space theory
We begin with the auxiliary space lemma [33, Lemma 3.1], which is closely related to the fictitious space method [29] and the auxiliary space method [40]. This lemma plays an important role in the analysis developed in this paper. The idea of the lemma can be traced back to [39], and it has since been widely used in the design and analysis of iterative methods; see, for example, [1, 7].
@@
Let $V$ and $\undertilde{V}$ be finite-dimensional vector spaces and let $\Pi\colon\undertilde{V}\to V$ be a surjective linear operator. Let $\undertilde{B}\colon\undertilde{V}\to\undertilde{V}$ be a SPD linear operator, and define
(3) $B=\Pi\undertilde{B}\Pi^{t}\colon V\to V.$
Then $B$ is SPD. Moreover, it satisfies
$(B^{-1}v,v)=\inf_{\undertilde{v}\in\undertilde{V},\ \Pi\undertilde{v}=v}(\undertilde{B}^{-1}\undertilde{v},\undertilde{v}),\quad v\in V.$
Now, let $A\colon V\to V$ and $B\colon V\to V$ be SPD linear operators, and consider the situation where $B$ is used as a preconditioner for the linear system associated with AA. In this case, the convergence rate of iterative methods can be estimated in terms of the condition number of the preconditioned operator BABA, given by
$\kappa(BA)=\frac{\lambda_{\max}(BA)}{\lambda_{\min}(BA)},$
@@
The following theorem, introduced in [33, Theorem 3.6] and proved using Lemma 2.1, provides sharp estimates for the extremal eigenvalues of BABA when the preconditioner $B$ is defined by (3) for some SPD linear operator $\undertilde{B}\colon\undertilde{V}\to\undertilde{V}$ and a surjective linear operator $\Pi\colon\undertilde{V}\to V$. Here, $\undertilde{V}$ is a finite-dimensional vector space, referred to as an auxiliary space. Thus, the preconditioner $B$ is constructed in terms of the operator $\undertilde{B}$ defined on the auxiliary space $\undertilde{V}$.
Theorem $2.$
Let $A\colon V\to V$ be a SPD linear operator, and let $B\colon V\to V$ be given in (3). Then we have
$\lambda_{\min}(BA)$ $=\left(\sup_{v\in V,\ \|v\|_{A}=1}\inf_{\undertilde{v}\in\undertilde{V},\ \Pi\undertilde{v}=v}(\undertilde{B}^{-1}\undertilde{v},\undertilde{v})\right)^{-1},$
$\lambda_{\max}(BA)$ $=\left(\inf_{v\in V,\ \|v\|_{A}=1}\inf_{\undertilde{v}\in\undertilde{V},\ \Pi\undertilde{v}=v}(\undertilde{B}^{-1}\undertilde{v},\undertilde{v})\right)^{-1}.$
The result in Theorem 2.2 can be extended to more general situations, for instance when AA is only semi-SPD, $B$ is a general linear operator, and $\undertilde{B}$ is semi-SPD. Further details on these extensions can be found in [33].
@@ 3.2 Elimination of the primal variable
From the saddle point system (1), one can eliminate the primal variable $u$ and obtain a reduced system formulated solely in terms of the dual variable pp. Assume first that AA is SPD. From the first equation in (1), we have
@@
In (7), $S$ is called the Schur complement. Because AA is SPD and $B$ is surjective, $S$ is of the form (3). In particular, by Lemma 2.1, we deduce that $S$ is SPD. Hence, the saddle point system (1) is equivalent to the SPD linear system (7) for $p$, followed by a back substitution to recover $u$.
The case when AA is only semi-SPD is more delicate, since (6) is not valid. Nevertheless, elimination of the primal variable $u$ is still possible when full information about the null space $\mathcal{N}(A)$ is available, that is, when an explicit basis of $\mathcal{N}(A)$ is known. Such a situation arises, for example, in Neumann boundary value problems for the Poisson equation and linear elasticity [11].
In the following, we summarize the arguments from [34, Section 2.2] and [37, Section 6.3] concerning the elimination of $u$. Let
@@
be an injective linear operator with range $\mathcal{N}(A)$. Such a map NN can be constructed as the matrix whose columns form a basis of $\mathcal{N}(A)$. From the first equation of (1), compatibility requires $f-B^{t}p\in\mathcal{R}(A)$, which is equivalent to
@@
for some $\xi\in\mathbb{R}^{\dim\mathcal{N}(A)}$, where $A^{+}\colon V\to V$ denotes a pseudoinverse of AA, i.e., $AA^{+}$ acts as the identity on $\mathcal{R}(A)$. Substituting (9) into the second equation of (1) yields
@@
Note that (10) reduces to (7) when AA is SPD.
Now, (8) and (10) form a coupled system in the variables $p$ and $\xi$. We next eliminate $\xi$. Since $\mathcal{R}(N)=\mathcal{N}(A)$, the condition (5) implies that BNBN is injective. Hence the orthogonal projection onto
@@
The computation of $P$ requires solving a linear system of $(BN)^{t}BN$, which can be carried out at negligible cost when $\dim\mathcal{N}(A)$ is small.
@@
(14) $S_{0}p_{0}=d_{0},\quad\text{where}\quad S_{0}=PSP^{t}\colon W_{0}\to W_{0},\ d_{0}=d-PSBN\eta\in W_{0}.$
Thus, we arrive at a linear system in the single unknown $p_{0}$, which represents the dual problem in the case when AA is semi-SPD. Proposition 3.1 shows that $S_{0}$ is SPD and independent of the particular choice of the pseudoinverse $A^{+}$ (cf. [34, Lemma 2.19]).
@@ 3.3 Iterative methods
Here, we discuss iterative methods for solving the saddle point problem (1). For simplicity, we assume that AA is SPD. Based on the surveys [2, 41], there are three major approaches in iterative methods.
The first approach is that we apply iterative methods for the SPD dual problems (7) and (14). In this case, we can utilize vast existing results on iterative methods for SPD linear systems. See [33] for a systematic approach for design and analysis of iterative methods for SPD linear systems.
@@
$\mathcal{L}=\begin{bmatrix}A^{-1}&0\\ 0&-S^{-1}\end{bmatrix}$
is optimal in the sense that the preconditioned operator $\mathcal{L}\mathcal{A}$ has only three distinct eigenvalues, namely 11 and $\tfrac{1}{2}(1\pm\sqrt{5})$. Thus, in this case, the central task is the construction of efficient preconditioners for AA and $S$ [2].
@@ 4 Estimates for the Schur complements
We begin with the case when AA is SPD. An obvious but important observation is that the Schur complement $S$ has the same structure as in (3). In Theorem 4.1, we establish sharp estimates for the extremal eigenvalues of $S$ defined in (7), obtained as direct consequences of Theorem $2.$
@@
In the saddle point problem (1), suppose that AA is SPD. Then the Schur complement $S$ given in (7) satisfies
@@
$V\leftarrow W,\quad\undertilde{V}\leftarrow V,\quad A\leftarrow I,\quad B\leftarrow S,\quad\Pi\leftarrow B,\quad\undertilde{B}\leftarrow A^{-1}.$
@@
$\lambda_{\max}(S)=\left(\inf_{0\neq q\in W}\inf_{v\in V,\ Bv=q}\frac{(Av,v)}{\|q\|^{2}}\right)^{-1}=\left(\inf_{0\neq v\in V}\frac{(Av,v)}{\|Bv\|^{2}}\right)^{-1}=\sup_{0\neq v\in V}\frac{\|Bv\|^{2}}{(Av,v)},$
@@
As a corollary of Theorem 4.1, we obtain Corollary 4.3, which appeared in [37, Lemma 9.1] (see also [3]).
@@
In the saddle point problem (1), suppose that AA is SPD. Then the Schur complement $S$ given in (7) satisfies
@@
where in the last inequality we use the fact that the minimizer $v\in V$ of $\|v\|^{2}/\|q\|^{2}$ subject to $Bv=q$ must lie in $\mathcal{N}(B)^{\perp}$. Meanwhile, we have
@@
Sometimes a preconditioner $L\colon W\to W$ is applied to the Schur complement $S$. Theorem 4.5 provides sharp estimates for the extremal eigenvalues of the preconditioned operator LSLS.
@@
In the saddle point problem (1), suppose that AA is SPD. Let $L\colon W\to W$ be a SPD linear operator. Then the Schur complement $S$ given in (7) satisfies
@@
$V\leftarrow W,\quad\undertilde{V}\leftarrow V,\quad A\leftarrow L,\quad B\leftarrow S,\quad\Pi\leftarrow B,\quad\undertilde{B}\leftarrow A^{-1}$
@@
An interesting special case arises when a right inverse of $B$ is explicitly available. In this situation, one can construct a special preconditioner $L$ using the right inverse of $B$ as a building block; examples include certain nonoverlapping domain decomposition methods, which we will discuss in Section 7. The following corollary, originally stated in [25, Lemma 2] with a rather involved proof, follows directly from Theorem 4.5.
@@
In the saddle point problem (1), suppose that AA is SPD. Assume further that a linear operator $\bar{B}\colon V\to W$ satisfies $B\bar{B}^{t}=I$, and define a preconditioner $L\colon W\to W$ by
@@
Next, we consider the case when AA is semi-SPD. Recall that the projected Schur complement $S_{0}$ given in (14) satisfies
$S_{0}=(PB)A^{+}(PB)^{t}.$
As discussed in the proof of Proposition 3.1, the operator $PB|_{\mathcal{R}(A)}$ is surjective and $\mathcal{R}((PB)^{t})\subset\mathcal{R}(A)$. Hence, in view of Proposition 3.1, we may regard $A^{+}$ and PBPB as
$A^{+}\colon\mathcal{R}(A)\to\mathcal{R}(A),\quad PB\colon\mathcal{R}(A)\to W_{0},$
where $W_{0}$ was defined in (11). This observation shows that the analysis of the semi-SPD case proceeds in essentially the same way as in the SPD case; see Theorems 4.9 and 4.10.
@@
In the saddle point problem (1), suppose that AA is semi-SPD. Then the projected Schur complement $S_{0}$ given in (14) satisfies
@@
In the saddle point problem (1), suppose that AA is semi-SPD. Let $L\colon W_{0}\to W_{0}$ be a SPD linear operator. Then the projected Schur complement $S_{0}$ given in (14) satisfies
@@
$V\leftarrow W_{0},\quad\undertilde{V}\leftarrow\mathcal{R}(A),\quad A\leftarrow L,\quad B\leftarrow S_{0},\quad\Pi\leftarrow PB,\quad\undertilde{B}\leftarrow A^{+}$
@@
In the saddle point problem (1), suppose that AA is semi-SPD. Assume further that a linear operator $\bar{B}\colon V\to W$ satisfies $B\bar{B}^{t}=I$, and define a preconditioner $L\colon W_{0}\to W_{0}$ by
@@
$\lambda_{\min}(LS_{0})\geq 1,\quad\lambda_{\max}(LS_{0})=\sup_{0\neq v\in\mathcal{R}(A)}\frac{(A(P\bar{B})^{t}P\bar{B}v,(P\bar{B})^{t}P\bar{B}v)}{(Av,v)}.$

### v2

@@ 1 Introduction
The purpose of this paper is to present sharp estimates for the extremal eigenvalues of the Schur complements arising in saddle point problems of the form (1). If $A$ is SPD, then the Schur complement $S=BA^{-1}B^{t}$ has the structure of an auxiliary space preconditioner [31, 42], where $A^{-1}$ corresponds to an iterative method defined on the auxiliary space $V$, and $B$ corresponds to the transfer operator mapping the auxiliary space $V$ to the original space $W$. A similar argument also applies in the case when $A$ is semi-SPD. Thus, the Schur complement $S$ can be analyzed by means of the recently proposed auxiliary space theory [35], which provides sharp estimates for iterative methods for both SPD and semi-SPD systems in a unified framework.
Estimates for the extremal eigenvalues of the Schur complement when $A$ is SPD are presented in Theorem 4.1, which are sharp in the sense that they are expressed as identities. Sharp estimates for the Schur complement preconditioned by a general SPD preconditioner are given in Theorem 4.5. These results refine existing abstract estimates for the Schur complement (see, e.g., [4, 27, 39]) in the sense that the earlier results can be recovered as straightforward corollaries. While our estimates follow directly from elementary linear algebra within the auxiliary space framework, the existing results required more involved arguments. Finally, our results extend naturally to the case when $A$ is semi-SPD, as shown in Theorems 4.9 and 4.10.
@@ 1.1 Notation
We summarize here the notation used throughout the paper. Let $V$ and $W$ be finite-dimensional vector spaces equipped with inner products $(\cdot,\cdot)$ and the associated norms $\|\cdot\|$. For any subspace $X\subset V$, we denote by $X^{\perp}$ its orthogonal complement. For a linear operator $L\colon V\to W$, the null space and the range are denoted by $\mathcal{N}(L)$ and $\mathcal{R}(L)$, respectively, and the adjoint of $L$ is denoted by $L^{t}\colon W\to V$.
@@ 2 Auxiliary space theory
We begin with the auxiliary space lemma [35, Lemma 3.1], which is closely related to the fictitious space method [31] and the auxiliary space method [42]. This lemma plays an important role in the analysis developed in this paper. The idea of the lemma can be traced back to [41], and it has since been widely used in the design and analysis of iterative methods; see, for example, [1, 8].
@@
Let $V$ and $\undertilde{V}$ be finite-dimensional vector spaces and let $\Pi\colon\undertilde{V}\to V$ be a surjective linear operator. Let $\undertilde{M}\colon\undertilde{V}\to\undertilde{V}$ be an SPD linear operator, and define
(3) $M=\Pi\undertilde{M}\Pi^{t}\colon V\to V.$
Then $M$ is SPD. Moreover, it satisfies
$(M^{-1}v,v)=\inf_{\undertilde{v}\in\undertilde{V},\ \Pi\undertilde{v}=v}(\undertilde{M}^{-1}\undertilde{v},\undertilde{v}),\quad v\in V.$
Now, let $L\colon V\to V$ and $M\colon V\to V$ be SPD linear operators, and consider the situation where $M$ is used as a preconditioner for the linear system associated with $L$. In this case, the convergence rate of iterative methods can be estimated in terms of the condition number of the preconditioned operator $ML$, given by
$\kappa(ML)=\frac{\lambda_{\max}(ML)}{\lambda_{\min}(ML)},$
@@
The following theorem, introduced in [35, Theorem 3.6] and proved using Lemma 2.1, provides sharp estimates for the extremal eigenvalues of $ML$ when the preconditioner $M$ is defined by (3) for some SPD linear operator $\undertilde{M}\colon\undertilde{V}\to\undertilde{V}$ and a surjective linear operator $\Pi\colon\undertilde{V}\to V$. Here, $\undertilde{V}$ is a finite-dimensional vector space, referred to as an auxiliary space. Thus, the preconditioner $M$ is constructed in terms of the operator $\undertilde{M}$ defined on the auxiliary space $\undertilde{V}$.
Theorem 2.2.
Let $L\colon V\to V$ be an SPD linear operator, and let $M\colon V\to V$ be given in (3). Then we have
$\lambda_{\min}(ML)$ $=\left(\sup_{v\in V,\ \|v\|_{L}=1}\inf_{\undertilde{v}\in\undertilde{V},\ \Pi\undertilde{v}=v}(\undertilde{M}^{-1}\undertilde{v},\undertilde{v})\right)^{-1},$
$\lambda_{\max}(ML)$ $=\left(\inf_{v\in V,\ \|v\|_{L}=1}\inf_{\undertilde{v}\in\undertilde{V},\ \Pi\undertilde{v}=v}(\undertilde{M}^{-1}\undertilde{v},\undertilde{v})\right)^{-1}.$
The result in Theorem 2.2 can be extended to more general situations, for instance when $L$ is only semi-SPD, $M$ is a general linear operator, and $\undertilde{M}$ is semi-SPD. Further details on these extensions can be found in [35].
@@ 3.2 Elimination of the primal variable
From the saddle point system (1), one can eliminate the primal variable $u$ and obtain a reduced system formulated solely in terms of the dual variable $p$. Assume first that $A$ is SPD. From the first equation in (1), we have
@@
In (7), $S$ is called the Schur complement. Because $A$ is SPD and $B$ is surjective, $S$ is of the form (3). In particular, by Lemma 2.1, we deduce that $S$ is SPD. Hence, the saddle point system (1) is equivalent to the SPD linear system (7) for $p$, followed by a back substitution to recover $u$.
The case when $A$ is only semi-SPD is more delicate, since (6) is not valid. Nevertheless, elimination of the primal variable $u$ is still possible when full information about the null space $\mathcal{N}(A)$ is available, that is, when an explicit basis of $\mathcal{N}(A)$ is known. Such a situation arises, for example, in Neumann boundary value problems for the Poisson equation and linear elasticity [12].
In the following, we summarize the arguments from [36, Section 2.2] and [39, Section 6.3] concerning the elimination of $u$. Let
@@
be an injective linear operator with range $\mathcal{N}(A)$. Such a map $N$ can be constructed as the matrix whose columns form a basis of $\mathcal{N}(A)$. From the first equation of (1), compatibility requires $f-B^{t}p\in\mathcal{R}(A)$, which is equivalent to
@@
for some $\xi\in\mathbb{R}^{\dim\mathcal{N}(A)}$, where $A^{+}\colon V\to V$ denotes a pseudoinverse of $A$, i.e., $AA^{+}$ acts as the identity on $\mathcal{R}(A)$. Substituting (9) into the second equation of (1) yields
@@
Note that (10) reduces to (7) when $A$ is SPD.
Now, (8) and (10) form a coupled system in the variables $p$ and $\xi$. We next eliminate $\xi$. Since $\mathcal{R}(N)=\mathcal{N}(A)$, the condition (5) implies that $BN$ is injective. Hence the orthogonal projection onto
@@
The computation of $P$ requires solving a linear system involving $(BN)^{t}BN$, which can be carried out at negligible cost when $\dim\mathcal{N}(A)$ is small.
@@
(14) $S_{0}p_{0}=d_{0},\quad\text{where}\quad S_{0}=PSP^{t}\colon W_{0}\to W_{0},\ d_{0}=P(d-SBN\eta)\in W_{0}.$
Thus, we arrive at a linear system in the single unknown $p_{0}$, which represents the dual problem in the case when $A$ is semi-SPD. Proposition 3.1 shows that $S_{0}$ is SPD and independent of the particular choice of the pseudoinverse $A^{+}$ (cf. [36, Lemma 2.19]).
@@ 3.3 Iterative methods
Here, we discuss iterative methods for solving the saddle point problem (1). For simplicity, we assume that $A$ is SPD. Based on the surveys [2, 43], there are three major approaches to iterative methods.
The first approach is to apply iterative methods for the SPD dual problem (7) ((14) in the semi-SPD case). In this case, we can utilize vast existing results on iterative methods for SPD linear systems. See [35] for a systematic approach to the design and analysis of iterative methods for SPD linear systems.
@@
$\mathcal{L}=\begin{bmatrix}A^{-1}&0\\ 0&S^{-1}\end{bmatrix}$
is optimal in the sense that the preconditioned operator $\mathcal{L}\mathcal{A}$ has only three distinct eigenvalues, namely $1$ and $\tfrac{1}{2}(1\pm\sqrt{5})$. Thus, in this case, the central task is the construction of efficient preconditioners for $A$ and $S$ [2].
@@ 4 Estimates for the Schur complements
We begin with the case when $A$ is SPD. An obvious but important observation is that the Schur complement $S$ has the same structure as in (3). In Theorem 4.1, we establish sharp estimates for the extremal eigenvalues of $S$ defined in (7), obtained as direct consequences of Theorem 2.2.
@@
In the saddle point problem (1), suppose that $A$ is SPD. Then the Schur complement $S$ given in (7) satisfies
@@
$V\leftarrow W,\quad\undertilde{V}\leftarrow V,\quad L\leftarrow I,\quad M\leftarrow S,\quad\Pi\leftarrow B,\quad\undertilde{M}\leftarrow A^{-1}.$
@@
$\lambda_{\max}(S)=\left(\inf_{0\neq q\in W}\inf_{v\in V,\ Bv=q}\frac{(Av,v)}{\|q\|^{2}}\right)^{-1}=\left(\inf_{v\in V,\ Bv\neq 0}\frac{(Av,v)}{\|Bv\|^{2}}\right)^{-1}=\sup_{0\neq v\in V}\frac{\|Bv\|^{2}}{(Av,v)},$
@@
As a corollary of Theorem 4.1, we obtain Corollary 4.3, which appeared in [39, Lemma 9.1] (see also [4]).
@@
In the saddle point problem (1), suppose that $A$ is SPD. Then the Schur complement $S$ given in (7) satisfies
@@
where in the last equality we use the fact that the minimizer $v\in V$ of $\|v\|^{2}/\|q\|^{2}$ subject to $Bv=q$ must lie in $\mathcal{N}(B)^{\perp}$. Meanwhile, we have
@@
Sometimes a preconditioner $L\colon W\to W$ is applied to the Schur complement $S$. Theorem 4.5 provides sharp estimates for the extremal eigenvalues of the preconditioned operator $LS$.
@@
In the saddle point problem (1), suppose that $A$ is SPD. Let $L\colon W\to W$ be an SPD linear operator. Then the Schur complement $S$ given in (7) satisfies
@@
$V\leftarrow W,\quad\undertilde{V}\leftarrow V,\quad L\leftarrow L,\quad M\leftarrow S,\quad\Pi\leftarrow B,\quad\undertilde{M}\leftarrow A^{-1}$
@@
An interesting special case arises when a right inverse of $B$ is explicitly available. In this situation, one can construct a special preconditioner $L$ using the right inverse of $B$ as a building block; examples include certain nonoverlapping domain decomposition methods, which we will discuss in Section 7. The following corollary, originally stated in [27, Lemma 2] with a rather involved proof, follows directly from Theorem 4.5.
@@
In the saddle point problem (1), suppose that $A$ is SPD. Assume further that a linear operator $\bar{B}\colon V\to W$ satisfies $B\bar{B}^{t}=I$, and define a preconditioner $L\colon W\to W$ by
@@
Next, we consider the case when $A$ is semi-SPD. Recall that the projected Schur complement $S_{0}$ given in (14) satisfies
$S_{0}=PBA^{+}(PB)^{t}.$
As discussed in the proof of Proposition 3.1, the operator $PB|_{\mathcal{R}(A)}$ is surjective and $\mathcal{R}((PB)^{t})\subset\mathcal{R}(A)$. Hence, in view of Proposition 3.1, we may regard $A^{+}$ and $PB$ as
(15) $A^{+}\colon\mathcal{R}(A)\to\mathcal{R}(A),\quad PB\colon\mathcal{R}(A)\to W_{0},$
where $W_{0}$ was defined in (11). This observation shows that the analysis of the semi-SPD case proceeds in essentially the same way as in the SPD case; see Theorems 4.9 and 4.10. Consequently, Theorems 4.9 and 4.10 do not depend on a particular choice of the pseudoinverse $A^{+}$ on the whole space $V$, but only on the canonical inverse of $A$ on $\mathcal{R}(A)$.
@@
In the saddle point problem (1), suppose that $A$ is semi-SPD. Then the projected Schur complement $S_{0}$ given in (14) satisfies
@@
In the saddle point problem (1), suppose that $A$ is semi-SPD. Let $L\colon W_{0}\to W_{0}$ be an SPD linear operator. Then the projected Schur complement $S_{0}$ given in (14) satisfies
@@
$V\leftarrow W_{0},\quad\undertilde{V}\leftarrow\mathcal{R}(A),\quad L\leftarrow L,\quad M\leftarrow S_{0},\quad\Pi\leftarrow PB,\quad\undertilde{M}\leftarrow A^{+}$
@@
In the saddle point problem (1), suppose that $A$ is semi-SPD. Assume further that a linear operator $\bar{B}\colon V\to W$ satisfies $B\bar{B}^{t}=I$, and define a preconditioner $L\colon W_{0}\to W_{0}$ by
@@
$\lambda_{\min}(LS_{0})\geq 1,\quad\lambda_{\max}(LS_{0})=\sup_{0\neq v\in\mathcal{R}(A)}\frac{(A(P\bar{B})^{t}PBv,(P\bar{B})^{t}PBv)}{(Av,v)}.$